# A Constrained Saddle Search Approach for Constructing Singular and Flexible Bar Frameworks

### v1

@@ 1 Introduction
A common and systematic approach to studying singularities in robotic systems is through the analysis of the Jacobian matrix derived from geometric constraints. To illustrate this, we consider a well-known example: the four-bar linkage in fig. 1, a simple yet widely studied singular and flexible bar framework (see e.g. [4]). This particular 2D bar framework consists of four vertices and four edges, with two pairs of edges having equal lengths. When the edges $AB$ and $AD$ do not align into a straight line (as illustrated in the right figures of fig. 1), the four-bar linkage is non-singular and the non-trivial deformations that keep the lengths of all edges form a 1-dimensional smooth manifold. The only degree of freedom in this motion can be characterized as $\theta_{1}$ in fig. 1 ($\theta_{2}$ is determined by $\theta_{1}$). However, when the edges $AB$ and $AD$ align into a straight line with $\theta_{1}=\theta_{2}=0$ (or $\pi$), the four-bar linkage becomes singular. The $4\times 8$ Jacobian matrix of the bar framework’s forward kinematics, also known as the rigidity matrix, loses rank. In fact, the Jacobian matrix is rank 3 and its null space has a two-dimensional subspace of infinitesimal flexes that are not infinitesimal rigid body motions. However, not all infinitesimal flexes in this subspace originate from nonlinear flexes that keep the lengths of all bars. The configuration space at this singular structure has two branches of 1-dimensional nonlinear flexes (these two flexes can be viewed as the two deformations in fig. 1 with small $\theta_{1}$). The singularity among four-bar linkages can be viewed as a branch point where some infinitesimal flexes cannot be realized by nonlinear flexes.
Let us analyze this singular bar framework from a different perspective by allowing the edge $CD$ to vary its length. Without the edge $CD$, we have an extra degree of freedom $\theta_{2}$ to deform the bar framework without changing the lengths of edges $AD,AB$ and $BC$. When viewing the squared length of edge $CD$ as a function of $\theta_{1}$ and $\theta_{2}$ in fig. 1, we can see that there are two saddle points at $\theta_{1}=\theta_{2}=0$ (or $\pi$). Both saddle points correspond to the singular structures where edges $AB$ and $AD$ are colinear. In fact, the two branches of nonlinear flexes correspond to the two branches in the level set at $\theta_{1}=\theta_{2}=0$. The singularity of the bar framework originates from the level set’s singularity: while the 2D non-trivial infinitesimal flexes are parameterized by $\theta_{1}$ and $\theta_{2}$, only the two infinitesimal flexes along $\theta_{1}=\pm\theta_{2}$ are realizable within the level set.
@@
Now let us generalize such a perspective of viewing singularity as saddle points on a constrained manifold to general flexible bar frameworks. We start with an under-constrained bar framework in $\mathbb{R}^{d}$ with $n$ vertices and $m$ edges ($nd>m+d(d+1)/2$). We assume that this given bar framework is non-singular and the configuration space of deformations that do not change the lengths of bars forms a $(nd-m)$-dimensional manifold. To avoid rigid body motions, we choose a pinning scheme so that the non-trivial nonlinear flexes form a $\left(nd-m-d(d+1)/2\right)$-dimensional manifold. Then, we allow a certain edge to vary its length and take the squared length of this free edge as the energy functional. By searching for index-$k$ saddle points of this energy functional on the constrained manifold where the lengths of the rest edges are fixed, we are guaranteed to find a singular and flexible bar framework. Without loss of generality, we choose the first edge as the free edge, and our constrained saddle search problem can be stated as follows:
[…]
@@
Our main result, theorem 3.1, establishes that under mild conditions, a non-degenerate index-$k$ saddle points corresponds to a singular and flexible bar framework. We also provide a constrained index-$k$ saddle search algorithm and reveal new singular structures that do not have the symmetry of the four-bar linkage. We highlight that this approach offers significant flexibility in designing singular structures, as it does not require any inherent symmetry, unlike the commonly studied singular structures. It is worth mentioning that optimizing $f_{1}(\vec{p})$ with constraints in eqs. 2 and 3 yields rigid bar frameworks with second-order rigidity [2]. These bar frameworks, though rigid, can also be viewed as singular since they have first-order non-trivial flexes.
@@ 2 Preliminiaries and notation
A simple counting argument, dating back to Maxwell [6], is commonly used to suggest whether a bar framework is flexible. Maxwell’s counting argument states that for a flexible bar framework, the number of edges should not exceed the number of degrees of freedom minus the number of rigid body motions, i.e. a flexible bar framework in $\mathbb{R}^{d}$ with $n$ vertices and $m$ edges should satisfy $m+d(d+1)/2<nd$. We categorize these bar frameworks under-constrained. It is also tempting to consider there are $nd-m-d(d+1)/2$ internal degrees of freedom for nonlinear flexes among under-constrained bar frameworks. However, such counting arguments do not always work. In fact, it only works when there is no singularity in the configuration space.
@@
A bar framework is flexible if the level set $\mathcal{M}(\vec{p})$ includes non-trivial deformations other than rigid body motions. A flexible bar framework is singular if the level set $\mathcal{M}(\vec{p})$ is not a $(nd-m)$-dimensional smooth manifold. A necessary condition for the singularity in bar frameworks is that $\nabla f_{i}(\vec{p})$ with $i=1,\dots,m$ are linearly dependent. In fact, we call the following Jacobian matrix
@@
Therefore, the level set $\vec{c}(\vec{x})=\vec{0}$ forms a $(d-m)$-dimensional smooth manifold. This linear independence condition is also known as the linear independence constraint qualification (LICQ) condition in constrained optimization [9]. The normal space of the level set $\vec{c}(\vec{x})=\vec{0}$ is defined as $\vec{N}(\vec{x})=\text{span}\{\nabla c_{1}(\vec{x}),\dots,\nabla c_{m}(\vec{x})\}$, and the tangent space is defined as the orthogonal complement $\vec{T}(\vec{x})=\vec{N}(\vec{x})^{\perp}$. We also define the projection matrix $\vec{P_{T}}(\vec{x})$ that projects vectors to the tangent space $\vec{T}(\vec{x})$
@@
Consider the case where $0<k<d-m$. A point $\vec{x^{*}}$ is called a non-degenerate index-$k$ saddle point of the energy functional $E(\vec{x})$ subject to $\vec{c}(\vec{x})=\vec{0}$ if the following conditions hold:
@@
has $k$ negative eigenvalues and $d-m-k$ positive eigenvalues in $\vec{T}(\vec{x^{*}})$ and the $m$-dimensional $\vec{N}(\vec{x^{*}})$ form the zero eigenspace of $\vec{\widehat{H}}(\vec{x^{*}})$.
It is worth mentioning that the Lagrange multiplier $\vec{\eta}(\vec{x^{*}})$ in eq. 9 is always unique when the LICQ condition eq. 7 holds at $\vec{x^{*}}$. The non-degeneracy of a saddle point $\vec{x^{*}}$ on the constrained manifold where $\vec{c}(\vec{x})=\vec{0}$ indicates the eigenvalues on the tangent space $\vec{T}(\vec{x^{*}})$ never vanish. The index $k$ of the saddle point $\vec{x^{*}}$ indicates that there are $k$ negative directions and $m-d-k$ positive directions on the tangent space $\vec{T}(\vec{x^{*}})$. Applying the Morse lemma, by a change of coordinates, the energy functional $E(\vec{x})$ on the constrained manifold where $\vec{c}(\vec{x})=\vec{0}$ can be written as a quadratic form with $k$ negative directions and $m-d-k$ positive directions. The Morse lemma (see e.g. [5]) is stated as:
@@
Using the local coordinate $\vec{z}$, we can see that $\vec{z}=\vec{0}$ is an algebraic singularity since the level set $\mathcal{L}_{\vec{z}}=\{\vec{z}\>|\>E(\vec{z})=E(\vec{x^{*}})\}$ is neither a single point nor a smooth manifold. At $\vec{z}=\vec{0}$, the Jacobian matrix of $E(\vec{z})$ is rank-deficient and not all vectors in the tangent space originate from the level set. However, in some neighborhood of $\vec{0}$ but removing the point $\vec{0}$, the level set $\mathcal{L}_{\vec{z}}$ forms a $(d-m-1)$-dimensional smooth manifold. Back to coordinate $\vec{x}$, at $\vec{x^{*}}$, the level set $\mathcal{L}_{\vec{x}}=\{\vec{x}\>|\>E(\vec{x})=E(\vec{x^{*}})\text{ and }\vec{c}(\vec{x})=\vec{0}\}$ is neither a single point nor a smooth manifold. The saddle point $\vec{x^{*}}$ is an algebraic singularity and there are tangent vectors at $\vec{x^{*}}$ that do not originate from smooth curves within $\mathcal{L}_{\vec{x}}$.
@@ 3 Constrained saddle search approach
Given the constrained saddle search problem in eqs. 1, 2 and 3, we are interested in the non-degenerate index-$k$ saddle points $\vec{p^{*}}\in\mathbb{R}^{nd}$ ($20<k<nd-m-d(d+1)/$ 2). The new bar framework corresponding to $\vec{p^{*}}$ has the same connection as the initial bar framework and remains under-constrained. In fact, this new bar framework is guaranteed to be singular and flexible.
@@
Theorem 3.1 is a direct application of the Morse lemma 2.2 since the singularity in the configuration space comes from the algebraic singularity of the level set at the non-degenerate saddle point. Our goal is to utilize theorem 3.1 to provide a numerical algorithm to find singular and flexible bar frameworks. We provide algorithm 1 for searching index-$k$ saddle points in a general form with an energy functional $E(\vec{x})$ and constraints $\vec{c}(\vec{x})=\vec{0}$, which is a combination of unconstrained saddle search (see e.g. [11]) and Newton’s method to project the search search direction back to the constraint set. However, unlike gradient descent which is guaranteed to find local minima, saddle search algorithms only have local convergence. By trying multiple initial conditions, one might be able to obtain a saddle point numerically.
@@
Examples of singular and flexible bar frameworks: We present two examples of 2D singular and flexible bar frameworks identified through our constrained saddle search approach. Let us explain in detail using the example in fig. 2(a). We begin with a non-singular heptagon consisting of 10 edges, where non-trivial nonlinear flexes forms a one-dimensional smooth manifold. To prevent rigid body motion in 2D, here we fix three coordinates, the $x,y$ coordinates of vertex $A$ and the $y$ coordinate of vertex $G$. Next, we set the edge $BF$ free and this additional degree of freedom allows us to search for index-1 saddle points of the energy functional, the squared length of edge $BF$. By searching for index-1 saddle point numerically, we obtain a bar framework in fig. 2(b), which does not have the four-bar linkage as a subgraph and has a two-dimensional non-trivial infinitesimal flexes. By testing the non-degeneracy and the LICQ condition, we find that the new bar framework is singular and only two infinitesimal flexes (up to scaling) that come from nonlinear flexes. These two infinitesimal flexes are found by a second-order stress test and plotted in fig. 2(b) and (d). We also find the nonlinear flexes corresponding to these two infinitesimal flexes, and plot one deformed states along the nonlinear flexes in fig. 2(c) and (e) [^1] Animations of these nonlinear flexes and more examples of singular and flexible frameworks are available at https://xuenanli.github.io/research/.. We provide another singular and flexible bar framework of this kind in fig. 3.
@@
Figure 2: (a) initial non-singular heptagon; (b)(d) final state with two special infinitesimal flexes; (c)(e) deformed states along the two directions in (b)(d).
@@
Figure 3: Another singular and flexible bar frameworks. Figures (a)-(d) follow the same instruction as fig. 2 (b)-(e).
@@ 4 Conclusions and future work
In this paper, we interpret singularities in flexible bar frameworks as non-degenerate saddle points on constrained manifolds and provide a numerical algorithm to obtain new singular and flexible frameworks. Future directions include: (1) allowing multiple edges to vary their lengths to uncover singular frameworks with higher-index; (2) applying it to periodic frameworks to identify singular lattice systems and design novel metamaterials (see e.g. [3]); and (3) generalizing it to broader geometric constraints to explore singular structures in complex robotic systems.

### v2

@@ 1 Introduction
A common and systematic approach to studying singularities in robotic systems is through the analysis of the Jacobian matrix derived from geometric constraints. To illustrate this, we consider a well-known example: the four-bar linkage in fig. 1, a simple yet widely studied singular and flexible bar framework (see e.g. [9]). This particular 2D bar framework consists of four vertices and four edges, with two pairs of opposite edges having equal lengths. When the edges $AB$ and $AD$ do not align into a straight line (as illustrated in the right figures of fig. 1), the four-bar linkage is non-singular and the non-trivial deformations that keep the lengths of all edges form a 1-dimensional smooth manifold. The only degree of freedom in this motion can be characterized as $\theta_{1}$ in fig. 1 ($\theta_{2}$ is determined by $\theta_{1}$). However, when the edges $AB$ and $AD$ align into a straight line with $\theta_{1}=\theta_{2}=0$ (or $\pi$), the four-bar linkage becomes singular. The $4\times 8$ Jacobian matrix of the bar framework’s forward kinematics, also known as the rigidity matrix, loses rank. In fact, the Jacobian matrix is rank 3 and its null space has a two-dimensional subspace of infinitesimal flexes that are not infinitesimal rigid body motions. However, not all infinitesimal flexes in this subspace originate from nonlinear flexes that keep the lengths of all bars. The configuration space at this singular structure has two branches of 1-dimensional nonlinear flexes (these two flexes can be viewed as the two deformations in fig. 1 with small $\theta_{1}$). The singularity among four-bar linkages can be viewed as a branch point where some infinitesimal flexes cannot be realized by nonlinear flexes.
Let us analyze this singular bar framework from a different perspective by allowing the edge $CD$ to vary its length. Without the edge $CD$, we have an extra degree of freedom $\theta_{2}$ to deform the bar framework without changing the lengths of edges $AD,AB$ and $BC$. When viewing the squared length of edge $CD$ as a function of $\theta_{1}$ and $\theta_{2}$ in fig. 1, we can see that there are two saddle points at $\theta_{1}=\theta_{2}=0$ (or $\pi$). Both saddle points correspond to the singular structures where edges $AB$ and $AD$ are colinear. In fact, the two branches of nonlinear flexes correspond to the two branches in the level set at $\theta_{1}=\theta_{2}=0$. The singularity of the bar framework originates from the level set’s singularity: though the 2-dimensional non-trivial infinitesimal flexes can be parameterized by $\theta_{1}$ and $\theta_{2}$, only the directions tangent to the level set along $\theta_{1}=\pm\theta_{2}$ correspond to realizable nonlinear flexes within the constraint set.
@@
Now let us generalize such a perspective of viewing singularity as saddle points on a constrained manifold to general flexible bar frameworks. We start with an under-constrained bar framework in $\mathbb{R}^{d}$ with $n$ vertices and $m$ edges ($nd>m+d(d+1)/2$). We assume that this given bar framework is non-singular and the configuration space of deformations that do not change the lengths of bars forms a (ndm)(nd-m)-dimensional manifold. To avoid rigid body motions, we choose a pinning scheme so that the non-trivial nonlinear flexes form a $\left(nd-m-d(d+1)/2\right)$-dimensional manifold. Then, we allow a certain edge to vary its length and take the squared length of this free edge as the energy functional. By searching for index-$k$ saddle points of this energy functional on the constrained manifold where the lengths of the rest edges are fixed, we are guaranteed to find a singular and flexible bar framework. Without loss of generality, we choose the first edge as the free edge, and our constrained saddle search problem can be stated as follows:
$\text{find index-}k\text{ saddles for }$ $f_{1}(\vec{p})=|\vec{p}_{1,1}-\vec{p}_{1,2}|^{2},$ (1)
@@
Our main result, theorem 3.1, establishes that under certain mild conditions, a non-degenerate index-$k$ saddle point corresponds to a singular and flexible bar framework. We also provide a constrained index-$k$ saddle search algorithm and reveal new singular structures that do not have the symmetry of the four-bar linkage. We highlight that this approach offers significant flexibility in designing singular structures, as it does not require any inherent symmetry, unlike the commonly studied singular structures. It is worth mentioning that instead of searching for saddle points, if we optimize (minimize or maximize) $f_{1}(\vec{p})$ under the constraints in eqs. 2 and 3, then we obtain rigid bar frameworks with second-order rigidity [7]. Although these bar frameworks are rigid, they can still be viewed as singular due to the existence of non-trivial first-order flexes that fail to integrate into nonlinear flexes.
@@ 2 Preliminiaries and notation
A simple counting argument, dating back to Maxwell [11], is commonly used to suggest whether a bar framework is flexible. Maxwell’s counting argument states that for a flexible bar framework, the number of edges should not exceed the number of degrees of freedom minus the number of rigid body motions, i.e. a flexible bar framework in $\mathbb{R}^{d}$ with $n$ vertices and $m$ edges should satisfy $m+d(d+1)/2<nd$. We categorize these bar frameworks under-constrained. It is also tempting to consider there are ndmd(d+1)/2nd-m-d(d+1)/2 internal degrees of freedom for nonlinear flexes among under-constrained bar frameworks. However, such counting arguments do not always work. In fact, it only works when there is no singularity in the configuration space.
@@
A bar framework is flexible if the level set $\mathcal{M}(\vec{p})$ includes non-trivial deformations other than rigid body motions. A flexible bar framework is singular if the level set $\mathcal{M}(\vec{p})$ is not a (ndm)(nd-m)-dimensional smooth manifold. A necessary condition for the singularity in bar frameworks is that $\nabla f_{i}(\vec{p})$ with $i=1,\dots,m$ are linearly dependent. In fact, we call the following Jacobian matrix
@@
Therefore, the level set $\vec{c}(\vec{x})=\vec{0}$ forms a (dm)(d-m)-dimensional smooth manifold. This linear independence condition is also known as the linear independence constraint qualification (LICQ) condition in constrained optimization [14]. The normal space of the level set $\vec{c}(\vec{x})=\vec{0}$ is defined as
$\vec{N}(\vec{x})=\text{span}\{\nabla c_{1}(\vec{x}),\dots,\nabla c_{m}(\vec{x})\}$
and the tangent space is defined as the orthogonal complement $\vec{T}(\vec{x})=\vec{N}(\vec{x})^{\perp}$. We also define the projection matrix $\vec{P_{T}}(\vec{x})$ that projects vectors to the tangent space $\vec{T}(\vec{x})$
@@
Consider the case where 0<k<dm0<k<d-m. A point $\vec{x^{*}}$ is called a non-degenerate index-$k$ saddle point of the energy functional $E(\vec{x})$ subject to $\vec{c}(\vec{x})=\vec{0}$ if the following conditions hold:
@@
has $k$ negative eigenvalues and dmkd-m-k positive eigenvalues in $\vec{T}(\vec{x^{*}})$ and the $m$-dimensional $\vec{N}(\vec{x^{*}})$ form the zero eigenspace of $\vec{\widehat{H}}(\vec{x^{*}})$.
It is worth mentioning that the Lagrange multiplier $\vec{\eta}(\vec{x^{*}})$ in eq. 9 is always unique when the LICQ condition eq. 7 holds at $\vec{x^{*}}$. The non-degeneracy of a saddle point $\vec{x^{*}}$ on the constrained manifold where $\vec{c}(\vec{x})=\vec{0}$ indicates the eigenvalues on the tangent space $\vec{T}(\vec{x^{*}})$ never vanish. The index $k$ of the saddle point $\vec{x^{*}}$ indicates that there are $k$ negative directions and mdkm-d-k positive directions on the tangent space $\vec{T}(\vec{x^{*}})$. Applying the Morse lemma, by a change of coordinates, the energy functional $E(\vec{x})$ on the constrained manifold where $\vec{c}(\vec{x})=\vec{0}$ can be written as a quadratic form with $k$ negative directions and mdkm-d-k positive directions. The Morse lemma (see e.g. [10]) is stated as:
@@
Using the local coordinate $\vec{z}$, we can see that $\vec{z}=\vec{0}$ is an algebraic singularity since the level set $\mathcal{L}_{\vec{z}}=\{\vec{z}\>|\>E(\vec{z})=E(\vec{x^{*}})\}$ is neither a single point nor a smooth manifold. At $\vec{z}=\vec{0}$, the Jacobian matrix of $E(\vec{z})$ is rank-deficient and not all vectors in the tangent space originate from the level set. However, in some neighborhood of $\vec{0}$ but removing the point $\vec{0}$, the level set $\mathcal{L}_{\vec{z}}$ forms a (dm1)(d-m-1)-dimensional smooth manifold. Back to coordinate $\vec{x}$, at $\vec{x^{*}}$, the level set $\mathcal{L}_{\vec{x}}=\{\vec{x}\>|\>E(\vec{x})=E(\vec{x^{*}})\text{ and }\vec{c}(\vec{x})=\vec{0}\}$ is neither a single point nor a smooth manifold. The saddle point $\vec{x^{*}}$ is an algebraic singularity and there are tangent vectors at $\vec{x^{*}}$ that do not originate from smooth curves within $\mathcal{L}_{\vec{x}}$.
@@ 3 Constrained saddle search approach
Given the constrained saddle search problem in eqs. 1, 2 and 3, we are interested in the non-degenerate index-$k$ saddle points $\vec{p^{*}}\in\mathbb{R}^{nd}$ (0<k<ndmd(d+1)/20<k<nd-m-d(d+1)/2). The new bar framework corresponding to $\vec{p^{*}}$ has the same connection as the initial bar framework and remains under-constrained. In fact, this new bar framework is guaranteed to be singular and flexible.
@@
Theorem 3.1 is a direct application of the Morse lemma 2.2 since the singularity in the configuration space comes from the algebraic singularity [^1] The term algebraic singularity here refers to a configuration where the dimension of the space of infinitesimal flexes exceeds that of the nearby constraint level sets. Readers who are interested in a detailed classification of algebraic singularity types associated with specific singular bar frameworks are referred to [8]. of the level set at the non-degenerate saddle point. Our goal is to utilize theorem 3.1 to provide a numerical algorithm to find singular and flexible bar frameworks. We provide algorithm 1 for searching index-$k$ saddle points in a general form with an energy functional $E(\vec{x})$ and constraints $\vec{c}(\vec{x})=\vec{0}$, which is a combination of unconstrained saddle search (see e.g. [16]) and Newton’s method to project the search search direction back to the constraint set. However, unlike gradient descent which is guaranteed to find local minima, saddle search algorithms only have local convergence. By trying multiple initial conditions, one might be able to obtain a saddle point numerically.
@@
Remark 3.2.
It is worth noting that numerical algebraic geometry tools, such as Bertini (see e.g. [1, 2]), have been successfully used to compute saddle points in kinematic systems by solving systems of polynomial equations, particularly those with isolated zero solutions. However, Bertini’s performance tends to deteriorate as the number of degrees of freedom increases. In our constrained saddle search problem, the number of variables is not small due to the combination of free vertices and Lagrange multipliers associated with the constraints, which makes Bertini less practical for this setting.
Examples of singular and flexible bar frameworks: We present two examples of 2D singular and flexible bar frameworks identified through our constrained saddle search approach in fig. 2. Let us explain in detail using the example in fig. 2(a)-(e). We begin with a non-singular heptagon consisting of 10 edges, where non-trivial nonlinear flexes forms a one-dimensional smooth manifold. To avoid rigid body motion in 2D, here we fix three coordinates, the $x,y$ coordinates of vertex $A$ and the $y$ coordinate of vertex $G$. Next, we set the edge $BF$ free and this additional degree of freedom allows us to search for index-1 saddle points of the energy functional, the squared length of edge $BF$. By searching for index-1 saddle point numerically, we obtain a bar framework in fig. 2(b), which does not have the four-bar linkage as a subgraph and has a two-dimensional non-trivial infinitesimal flexes. By testing the non-degeneracy and the LICQ condition, we find that the new bar framework is singular and only two infinitesimal flexes (up to scaling) that come from nonlinear flexes. These two infinitesimal flexes are found by a second-order stress test and plotted in fig. 2(b) and (d). We also find the nonlinear flexes corresponding to these two infinitesimal flexes, and plot one deformed states along the nonlinear flexes in fig. 2(c) and (e) [^2] Animations of these nonlinear flexes and more examples of singular and flexible frameworks are available at https://xuenanli.github.io/research/..
@@
Figure 2: (a) initial non-singular heptagon; (b)(d) final state with two special infinitesimal flexes; (c)(e) deformed states along the two directions in (b)(d). Another singular and flexible bar framework is shown in figures (f)-(i), following the same instructions as those for figures (b)-(e).
@@ 4 Conclusions and future work
In this paper, we interpret singularities in flexible bar frameworks as non-degenerate saddle points on constrained manifolds and provide a numerical algorithm to obtain new singular and flexible frameworks. Future directions include: (1) exploring 3D singular bar frameworks; (2) allowing multiple edges to vary their lengths to uncover singular frameworks with higher-index; (3) generalizing our approach to periodic frameworks for identifying singular lattice systems and designing novel metamaterials (see e.g. [4, 5, 6]); and (4) generalizing it to broader geometric constraints to explore singular structures in complex robotic systems.